OpenStudy (anonymous):

PLEASE HELP ME, CHECK MY WORK/GIVE INPUT, I THINK I GOT THE ANSWER CORRECT :) Two different plants grow each year at different rates, which are represented by the functions f(x) = 4^x and g(x) = 5x + 2. What is the first year the f(x) height is greater than the g(x) height? (^x is an exponent)

OpenStudy (anonymous):

Year 0<<MY ANSWER Year 1 Year 2 Year 3 4^x=f(X) and g(x)=5x+2. 4^0= 0, right? so g(x)=2 0=f(x) and g(x)=2,

OpenStudy (anonymous):

well that is incorrect, read the question

OpenStudy (anonymous):

What is the first year the f(x) height is greater than the g(x) height? Is F(0) > G(0) in your equation?